The length of a rectangle room is 6 feet longer than twice the width, If the room's perimeter is 180 feet, what are the room's dimensions?
Thank You!

Answer by rfer(16322)

2x+2(2x+6)=180
2x+4x+12=180
6x=180-12
x=168/6
x=28 ft
2x+6=62 ft